) If F(x, y, z) = e xˆi + yzˆj − yz2ˆk, find the divergence of F at (0, 2, −1)


1
Expert's answer
2021-11-04T19:58:11-0400

If "F(x, y, z) = e^x\cdot i + yz\cdot j − yz^2\cdot k", let us find the divergence of "F" at "(0, 2, \u22121)."

It follows that

"Div\ [F(x,y,z)]=\frac{\partial( e^x)}{\partial x}
+\frac{\partial(yz)}{\partial y}
+\frac{\partial(-yz^2)}{\partial z}
=e^x+z-2yz,"

and hence

"Div\ [F(0,2,-1)]=e^0-1-2\cdot 2(-1)=4."


Answer: the divergence of "F" at "(0, 2, \u22121)" is equal to 4.